OpenStudy (anonymous):

write the equation in standard form, y^2 +4x-6y+5=0

OpenStudy (jdoe0001):

hmm do you know what a "perfect square trinomial" is?

OpenStudy (anonymous):

1) (y-5)(y-1) = 0 2) (x+7)^2 = 0 3) 5(x^2 - 25) = 0 (x-5)(x+5) = 0 x = 5 or x = -5

OpenStudy (anonymous):

yes sortof but its been a while since then

OpenStudy (anonymous):

Whenever you multiply a binomial by itself twice, the resulting tri-nominal is called a perfect square tri-nominal An Example: (x + 1) × (x + 1) = x2 + x + x + 1 = x2 + 2x + 1 and x2 + 2x + 1 is a perfect square tri-nominal

OpenStudy (anonymous):

i cant go to links my mom has it set up so i have to email her for permission, im trying to get the equation in standard form so i can get the vertex, focus, and the equation of the directrix of the porabola

OpenStudy (anonymous):

at least i think its a parabola

OpenStudy (jdoe0001):

ahemm well... to do that you'd need to check your what a perfect square trinomial is first

OpenStudy (anonymous):

ok i looked at it

OpenStudy (jdoe0001):

..... so...do you recall what it's?

OpenStudy (jdoe0001):

hmmm that'd be the difference of cubes

OpenStudy (anonymous):

ok then i have no idea :'(

OpenStudy (jdoe0001):

http://math.tutorvista.com/algebra/perfect-square-trinomial.html <--- that should refresh it.... try to see if you can get to the link

OpenStudy (anonymous):

my mom said no. uggghhhh why wont she be nice